OpenStudy (anonymous):

Find the derivative of int {3x}^{9x} of {u+6}/{u-1}du

OpenStudy (amistre64):

the derivative of in integration is f(b) b' - f(a) a' i think

OpenStudy (amistre64):

\[\frac{d}{dx}\int_{a}^{b}f'(u)du=\frac{d}{dx}F(b)-F(a)=f'(b)b'-f'(a)a'\]

OpenStudy (amistre64):

in other words, they already gave you f'

OpenStudy (anonymous):

= [ 9 (9x + 6 )/ ( 9x -1) ] - [ 3 ( 3x + 6) / ( 3x -1) ]